# Critical exponents of the Ising model with quenched structural disorder and long-range interactions at spatial dimension $d=3$

### v2

@@ 1 Introduction
This year marks the hundredth anniversary of Ernst Ising’s paper [1], which has introduced the famous now Ising model. The model was suggested to E. Ising by his adviser Wilhelm Lenz in order to describe a ferromagnetic phase transition. E. Ising has found no ferromagnetic ordering in one-dimensional chain of scalar magnetic moments (spins) with interaction between nearest neighbours for finite temperature [1]. However since then, the generalization of the Ising model for higher dimensions and its modifications have been applied to explain collective behaviour in various systems from a wide range of scientific fields both within and beyond the natural sciences [2]. The critical properties of the Ising model were also in the focus of interests of the brilliant Ukrainian physicist Ihor Yukhnovskii, whose hundredth anniversary is also celebrated this year. Especially he was interested in the effects of interaction range on the cooperative behaviour [3]. Authors of this paper studied a lot from his books [3, 4]. Therefore, it is a great honor for us to dedicate our paper on the Ising model to the memory of Prof. I. Yukhnovskii.
Our study concerns the analysis changes of critical behaviour of Ising model caused by mutual effects of two factors: long-range (LR) interactions and quenched disorder. The influence of a range of microscopic interactions between system components on critical behaviour is one of the most important problems in the theory of critical phenomena. Study of LR interactions are highly significant due to their prevalence in physical, chemical, and biological systems [5, 6, 7, 8] . Such interactions are usually modelled as algebraically decayed with distance $x$ between interacting particles according to power law
@@
where $d$ stands for space dimension. Models with interactions given by (1.1) have found wide-ranging applications in the different fields of physics [5, 6, 7, 8]. Recent advances in technology for realizing controllable LR interactions in optical lattices, molecular and atomic systems [9, 10], particularly with the possibility of continuous tuning of the decay parameter $\sigma$ [11], has further increased interest in the theoretical study of systems with interactions of type (1.1).
@@
Models with LR interactions posses distinctive features that affect their critical behaviour. Further on, we focus on the case of weak LR interactions, for which the decay parameter is positive, $\sigma>0$. For this case, the one-dimensional Ising model ($d=1$) with $\sigma<1$ in contrast to its short-range (SR) counterpart, was proven to exhibit phase transition to the long-range-ordered phase at non-zero temperature [12]. Field-theoretical renormalization group (RG) analysis of the LR interacting $n$-vector model (Ising system is its particular case at $n=1$) reveals that there are three universality classes depending on $\sigma$: (i) for $\sigma\leq d/2$ the classical mean-field critical behaviour is realised, (ii) for $\sigma\geq 2$ critical exponents coincides with those of the model with SR interactions, (iii) for $d/2<\sigma<2$ there is the LR universality class, for which critical exponents depend on $\sigma$ [13]. Later it was established that the boundary $\sigma^{*}$ between the SR and the LR universality classes is actually defined by pair correlation function critical exponent of the SR model $\eta_{SR}$, $\sigma^{*}=2-\eta_{SR}$ [14]. Such picture was corroborated by other approaches including non-perturbative RG (NPRG) [15], Monte Carlo simulations [16] and conformal bootstrap [17] (for other references see [18]). The estimates characterising the LR universality class have been obtained within Monte Carlo simulations for the Ising model mainly in spatial dimensions $d=1$ and $d=2$ (for collection of references see [19]). Only a few Monte Carlo results are available for $d=3$ [20, 21]. RG results for these exponents are available in low-loop approximations [13, 22, 23]. The study within the highest accessible three-loop approximation was performed only recently [19] with correction of loop-integral calculations in this year [24].
@@
Another factor discussed in this paper is the impact of structural disorder on the critical behaviour. Since some degree of structural imperfection is unavoidably present in almost all materials, this issue still attracts interest of researchers. It is known that the changes introduced by the disorder depend on its type: various realisations of quenched disorder have been studied for SR case (see e.g. [25]). However only two types of disorder have been considered for LR interacting systems in the form of dilution [26, 27, 28, 29, 30, 31, 32, 33] and in the form of random fields [34, 35, 36, 37, 38, 39].
@@
In its turn, LR interacting models with random field have been investigated by a bunch of approaches. In particular, different aspects of long-range low-temperature order have been explored by rigorous [34, 35, 36] as well as RG methods [40, 37]. Dynamical ordering has been analysed numerically by Monte Carlo simulations [38]. The question of possible dimensional reduction has been addressed within NPRG approach [39].
@@
In contrast, the diluted case has been studied only by field-theoretical RG methods [26, 27, 28, 29, 30, 31, 32, 33]. These studies demonstrated relevance of quenched dilution for critical behaviour and emergence of disorder-induced universality class. Because of the degeneracy of RG functions for the Ising model with uncorrelated disorder on the one-loop level the universal characteristics of this universality class have been obtained in the form of $\sqrt{\epsilon}$-expansions [28, 30, 31, 33]. These expansions are known to have bad convergence properties in the context of the study of SR models (see e. g. [41]). There has been only one attempt to get reliable results for fixed space dimension in this case: resummation of the RG functions was performed within a massive renormalization scheme [29]. Almost all previous estimates have been obtained within two-loop approximation only, while $\sqrt{\epsilon}$-expansion extracted from three-loop approximation [33] was based on the RG data with errors [19], corrected only recently [24]. Therefore, in our study we complete existing data by our estimate of correlation length critical exponent based on resummation of three-loop RG functions.
@@
The structure of the paper is as follows. In the next Section 2, we introduce the model and its field-theoretical formulation. In Section 3, we describe the renormalization procedure used and present RG functions. We treat these functions in the Section 4 to extract values of the critical exponent $\nu$ at spatial dimension $d=3$ with help of resummation procedures. In Section 5 we summarize our findings with conclusions.
@@ 2 Model
Figure 1: Structurally disordered two-dimensional ($d=2$) diluted Ising model: yellow disks represent lattice sites $\bf x$} occupied by Ising spins ${S}_{\bf x}$ (blue arrows), while dark gray sites correspond to non-magnetic impurities.
@@
${\cal H}=-\sum_{{\bf x},{\bf x}^{\prime}}J(|{\bf x}-{\bf x}^{\prime}|)c_{{\bf x}}c_{{\bf x}^{\prime}}{S}_{\bf x}{S}_{{\bf x}^{\prime}}\,$ (2.1)
where $\bf x$} runs over site of $d$-dimensional (hyper)cubic lattice, $J(|{\bf x}-{\bf x}^{\prime}|)>0$ is the ferromagnetic interaction between $S=\pm 1$ “classical” Ising spins on sites $\bf x$} and $\bf x$$^{\prime}$, its form is defined by Eq. (1.1), and the occupation numbers $c_{{\bf x}}$ either equal 1 if the lattice site $\bf x$} is occupied by a spin or equal 0 if site $\bf x$} is vacant. Here, we are interested in the case of the so-called weak dilution, when concentration of occupied sites $c$ is much larger than percolation threshold. We consider that all $c_{\bf x}$ are i.i.d. random variables distributed according to
${\cal P}(\{c_{{\bf x}}\})=\prod_{\bf x}p(c_{{\bf x}}),\hskip 28.45274ptp(c_{{\bf x}})=c\delta(c_{{\bf x}}-1)+(1-c)\delta(c_{{\bf x}}).$ (2.2)
We are interested in the case of quenched disorder, where the positions of vacancies/nonmagnetic impurities $c_{\bf x}$ are fixed (see Fig. 1). The static and dynamic critical behaviour of such random-site Ising model with interactions between neighbour spins only has been studied by many Monte Carlo simulations (see e.g., reviews in [42, 41]). Unfortunately, we do not know of any studies of such model when interactions decay with distance according to Eq. (1.1).
For systems with quenched disorder thermodynamics of physical observables can be analysed from the configurational average of the free energy [43]:
@@
$\overline{\left.\dots\right.}=\prod_{\bf x}\sum_{c_{\bf x}}p(c_{\bf x})(\dots)$ (2.5)
To avoid averaging of the logarithm in Eq. (2.3) one can utilize the replica trick [44, 45, 46]
$\ln Z(\{c\})=\lim_{m\to 0}\frac{\big(Z(\{c\})\big)^{m}-1}{m}\,.$ (2.6)
The $m$th power of the configuration-dependent partition function is represented in the form of a functional integral using the Stratonovich-Hubbard transformation. This transformation introduces field variables $\vec{\phi}^{\alpha}_{{\bf x}}$, ($\alpha=\{1,\dots,m\}$), conjugated to the spin variables. Subsequently, the trace over the spin subsystem is taken, and after configurational averaging, the free energy is obtained (see e.g. [47]):
${F}\sim\int({\rm d}\vec{\phi})\mathrm{e}^{-{\cal H}_{\rm eff}},\qquad\int({\rm d}{\phi})=\prod_{\alpha=1}^{n}\prod_{\bf{x}}\int_{-\infty}^{\infty}{\rm d}{\phi}^{\alpha}_{\bf x}\,$, (2.7)
@@
${\mathcal{H}}_{eff}{=}\int d^{d}x\left\{\frac{1}{2}[\tau|{\phi}^{\alpha}|^{2}{+}|\nabla^{\sigma/2}{\phi}^{\alpha}|^{2}]{+}\frac{u_{0}}{4!}\sum_{\alpha=1}^{m}|{\phi}^{\alpha}|^{4}{+}\frac{v_{0}}{4!}\left(\sum_{\alpha=1}^{m}({\phi}^{\alpha})^{2}\right)^{2}\right\}\,.$ (2.8)
In the limit $m\rightarrow 0$ field theory (2.8) describes the critical properties of a LR interacting random Ising model. The parameter $\tau$ is connected with the temperature distance to the critical point. The bare coupling $u_{0}$ is proportional to the concentration $\sim c$ and is positive whereas the bare coupling $v_{0}$ is proportional to $\sim c(c-1)$ and thus is negative. The last term in (2.8) is present only for non-zero dilution: it is directly responsible for the fluctuations effective interaction due to the presence of impurities. We remind that (2.8) was obtained with assumption of the weak quenched dilution (far above the percolation threshold). Note that theory (2.8) is valid only for $\sigma<2-\eta_{SR}$. Since we are interested only in the long-distance properties of the theory, in the limit of small wave vector $k\to 0$, we kept only first non-trivial term of the expansion of the Fourier image of interactions $k^{\sigma}$.
@@ 3 RG approach
Within the field-theoretical RG approach [48, 49, 50], the analysis of the critical properties of $\phi^{4}$ models is based on the calculation of so-called one-particle irreducible vertex functions. These functions are perturbatively calculated around the Gaussian model ($u=v=0$ in Eq. (2.8)) in the coupling constants $u$ and $v$. However, these calculations result in integrals that are divergent at the upper critical dimension, which in our case is $d_{u}=2\sigma$. To remove these divergencies, a controlled rearrangement of the perturbative series for the vertex functions is performed by imposing certain normalizing conditions. In the following, we use the massless minimal subtraction scheme [51, 52]. We express bare parameters via the renormalized ones, introducing the renormalizing factors of the field $\phi$, $Z_{\phi}$, the couplings, $Z_{u}$ and $Z_{v}$ and for operator $\phi^{2}$, $Z_{\phi^{2}}$ ensuring finiteness of two-point vertex function with one $\phi^{2}$ insertion.
Then, the scaling bahaviour can be analysed from RG $\beta$- and $\gamma$-functions characterising change of the vertex functions under the RG transformation. They are introduced as follows:
$\beta_{u}=\frac{\partial u}{\partial\ln{\mu}},\,\,\beta_{v}=\frac{\partial v}{\partial\ln{\mu}},\,\,\gamma_{\phi}=\frac{\partial Z_{\phi}}{\partial\ln{\mu}},\,\,\overline{\gamma}_{\phi^{2}}=-\frac{\partial\overline{Z}_{\phi^{2}}}{\partial\ln{\mu}},$ (3.1)
where $\overline{Z}_{\phi^{2}}=Z_{\phi^{2}}Z_{\phi}$ and the derivatives are taken at fixed bare parameters. Here $\mu$} is the scale parameter introduced in the same way as in [19] .
@@
The stability of the FPs (3.2) is defined by the stability matrix $\left(\begin{matrix}\frac{\partial\beta_{{u}}}{\partial u}&\frac{\partial\beta_{{u}}}{\partial v}\\ \frac{\partial\beta_{{v}}}{\partial u}&\frac{\partial\beta_{{v}}}{\partial v}\end{matrix}\right)\,.$ If this matrix, evaluated at a FP $(u^{*},v^{*})$, has both its eigenvalues $\lambda_{1}$ and $\lambda_{2}$ with positive real parts, then this FP is stable.
@@
All other critical exponents can be calculated from scaling relations [48, 49, 50].
We work in the three-loop approximation, using the results of loop integral calculations [19, 24]. Note that our procedure differs. The authors of Refs. [19, 24] fixed spatial dimension $d$ and looked for poles in the deviation of the decay parameter from the real space dimension $\sigma=(d-\epsilon)/2$. While we fix $\sigma$ and look for poles in the deviations of spatial dimension from the upper critical dimension $d=2\sigma-\epsilon$. The result of our calculations of the RG functions (3.1) within the three-loop approximation is as follows:
$\beta_{u}$ $=$ $u\Big(-\epsilon+\frac{3u}{2}+2v-\frac{1}{18}\left(27u^{2}+72uv+42v^{2}\right)D_{\sigma}$ (3.4)
@@
$+\frac{1}{216}(3u+2v)\left(9u^{2}+30uv+28v^{2}\right)H_{\sigma}\Big),$
$\beta_{v}$ $=$ $v\Big(-\epsilon+6u+\frac{4}{3}v-\frac{1}{24}\left(9u^{2}+36uv+22v^{2}\right)D_{\sigma}+$ (3.5)
$\frac{1}{24}\left(21u^{3}+94u^{2}v+129uv^{2}+48v^{3}\right)D_{\sigma}^{2}+\frac{1}{108}v\left(63u^{2}+111uv+44v^{2}\right)E_{\sigma}+$
$\frac{1}{{72}}\left(9u^{3}+30u^{2}v+45uv^{2}+16v^{3}\right)F_{\sigma}+\frac{1}{108}(3u+4v)\left(3u^{2}+6uv+2v^{2}\right)G_{\sigma}+$
$\frac{1}{108}v\left(9u^{2}+24uv+11v^{2}\right)H_{\sigma}\Big),$
$\bar{\gamma}_{\phi^{2}}$ $=$ $\frac{1}{2}u+\frac{1}{3}v-\frac{1}{12}\left(3u^{2}+6uv+2v^{2}\right)D_{\sigma}+$ (3.6)
$\frac{1}{864}\left(378u^{3}+1125u^{2}v+990uv^{2}+220v^{3}\right)D_{\sigma}^{2}+\frac{1}{432}v\left(9u^{2}+18uv+4v^{2}\right)E_{\sigma}+$
$\frac{1}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}96}}\left(6u^{3}+19u^{2}v+18uv^{2}+4v^{3}\right)F_{\sigma}+\frac{1}{216}(3u+2v)\left(3u^{2}+6uv+2v^{2}\right)G_{\sigma},$

### v3

@@ 1 Introduction
This year marks the centenary of Ernst Ising’s paper [1], which has introduced the nowadays famous Ising model. The model was suggested to Ernst Ising by his adviser Wilhelm Lenz in order to describe a ferromagnetic phase transition. Ernst Ising found no ferromagnetic ordering in one-dimensional chain of scalar magnetic moments (spins) with interaction between nearest neighbours for finite temperature [1]. However, since then, the generalization of the Ising model for higher dimensions and its modifications has been used to explain a collective behaviour in various systems from a wide range of scientific fields both within and beyond the natural sciences [2]. The critical properties of the Ising model were also in the focus of interests of the brilliant Ukrainian physicist Ihor Yukhnovskii, whose hundredth anniversary is also celebrated this year. Especially he was interested in the effects of the interaction range on the cooperative behaviour [3]. Authors of this paper studied a lot from his books [3, 4]. Therefore, it is a great honor for us to dedicate our paper on the Ising model to the memory of Prof. Ihor Yukhnovskii.
Our study concerns the analysis of the changes of critical behaviour of Ising model caused by mutual effects of two factors: long-range (LR) interactions and quenched disorder. The influence of a range of microscopic interactions between system components on critical behaviour is one of the most important problems in the theory of critical phenomena. Studies of LR interactions are highly significant due to their prevalence in physical, chemical, and biological systems [5, 6, 7, 8]. Such interactions are usually modelled as algebraically decayed with distance $x$ between interacting particles according to the power law
@@
where $d$ stands for space dimension. Models with interactions given by (1.1) have found wide applications in different fields of physics [5, 6, 7, 8]. Recent advances in technology for realizing controllable LR interactions in optical lattices, molecular and atomic systems [9, 10], particularly with the possibility of continuous tuning of the decay parameter $\sigma$ [11], have further increased interest in the theoretical study of systems with interactions of type (1.1).
@@
Models with LR interactions possess distinctive features that affect their critical behaviour. Further on, we focus on the case of weak LR interactions, for which the decay parameter is positive, $\sigma>0$. For this case, the one-dimensional Ising model ($d=1$) with $\sigma<1$ in contrast to its short-range (SR) counterpart, was proven to exhibit a phase transition to the long-range-ordered phase at non-zero temperature [12]. Field-theoretical renormalization group (RG) analysis of the LR interacting $n$-vector model (Ising system is its particular case at $n=1$) reveals that there are three universality classes depending on $\sigma$: (i) for $\sigma\leqslant d/2$ the classical mean-field critical behaviour is realised, (ii) for $\sigma\geqslant 2$ critical exponents coincides with those of the model with SR interactions, (iii) for $d/2<\sigma<2$ there is the LR universality class, for which critical exponents depend on $\sigma$ [13]. Later it was established that the boundary $\sigma^{*}$ between the SR and the LR universality classes is actually defined by a pair correlation function critical exponent of the SR model $\eta_{\rm SR}$, $\sigma^{*}=2-\eta_{\rm SR}$ [14]. Such a picture was corroborated by other approaches including non-perturbative RG (NPRG) [15], Monte Carlo simulations [16] and conformal bootstrap [17] (for other references see [18]). The estimates characterizing the LR universality class have been obtained within Monte Carlo simulations for the Ising model mainly in spatial dimensions $d=1$ and $d=2$ (for collection of references see [19]). Only a few Monte Carlo results are available for $d=3$ [20, 21]. RG results for these exponents are available in low-loop approximations [13, 22, 23]. The study within the highest accessible three-loop approximation was performed only recently [19] with a correction of loop-integral calculations this year [24].
@@
Another factor discussed in this paper is the impact of structural disorder on the critical behaviour. Since some degree of structural imperfection is unavoidably present in almost all materials, this issue still attracts interest of researchers. It is known that the changes introduced by the disorder depend on its type: various realizations of quenched disorder have been studied for SR case (see, e.g., [25]). However, only two types of disorder have been considered for LR interacting systems in the form of dilution [26, 27, 28, 29, 30, 31, 32, 33] and in the form of random fields [34, 35, 36, 37, 38, 39].
@@
In its turn, LR interacting models with random field have been investigated by a bunch of approaches. In particular, different aspects of low-temperature long-range order have been explored by rigorous [34, 35, 36] as well as RG methods [40, 37]. Dynamical ordering has been analysed numerically by Monte Carlo simulations [38]. The issue of possible dimensional reduction has been addressed within NPRG approach [39].
@@
By contrast, the diluted case has been studied only by field-theoretical RG methods [26, 27, 28, 29, 30, 31, 32, 33]. These studies demonstrated relevance of quenched dilution for critical behaviour and the emergence of disorder-induced universality class. Due to the degeneracy of RG functions for the Ising model with uncorrelated disorder at the one-loop level, the universal characteristics of this universality class have been obtained in the form of $\sqrt{\epsilon}$-expansions [28, 30, 31, 33]. These expansions are known to have bad convergence properties in the context of the study of SR models (see, e.g., [41]). There has been only one attempt to get reliable results for fixed space dimension in this case: resummation of the RG functions was performed within a massive renormalization scheme [29]. Almost all previous estimates have been obtained within the two-loop approximation only, while $\sqrt{\epsilon}$-expansion extracted from three-loop approximation [33] was based on the RG data with errors [19], corrected only recently [24]. Therefore, in our study we complete the existing data by our estimate of correlation length critical exponent based on resummation of three-loop RG functions.
@@
The structure of the paper is as follows. In the next section 2, we introduce the model and its field-theoretical formulation. In section 3, we describe the renormalization procedure used and present RG functions. We treat these functions in section 4 to extract values of the critical exponent $\nu$ at spatial dimension $d=3$ with the help of resummation procedures. In section 5 we summarize our findings with conclusions.
@@ 2 Model
Figure 1: (Colour online) Structurally disordered two-dimensional ($d=2$) diluted Ising model: yellow disks represent lattice sites $\bf x$} occupied by Ising spins ${S}_{\bf x}$ (blue arrows), while dark gray sites correspond to non-magnetic impurities.
@@
${\cal H}=-\sum_{{\bf x},{\bf x}^{\prime}}J(|{\bf x}-{\bf x}^{\prime}|)c_{{\bf x}}c_{{\bf x}^{\prime}}{S}_{\bf x}{S}_{{\bf x}^{\prime}},$ (2.1)
where $\bf x$} runs over site of $d$-dimensional (hyper)cubic lattice, $J(|{\bf x}-{\bf x}^{\prime}|)>0$ is the ferromagnetic interaction between $S=\pm 1$ “classical” Ising spins on sites $\bf x$} and $\bf x$$^{\prime}$, its form is defined by equation (1.1), and the occupation numbers $c_{{\bf x}}$ either equal to 1 if the lattice site $\bf x$} is occupied by a spin or equal to 0 if site $\bf x$} is vacant. Here, we are interested in the case of the so-called weak dilution, when concentration of the occupied sites $c$ is much larger than percolation threshold. We consider that all $c_{\bf x}$ are independent and identically distributed (i.i.d.) random variables distributed according to
${\cal P}(\{c_{{\bf x}}\})=\prod_{\bf x}p(c_{{\bf x}}),\qquad p(c_{{\bf x}})=c\delta(c_{{\bf x}}-1)+(1-c)\delta(c_{{\bf x}}).$ (2.2)
We are interested in the case of quenched disorder, where the positions of vacancies/nonmagnetic impurities $c_{\bf x}$ are fixed (see figure 1). The static and dynamic critical behaviour of such random-site Ising model with interactions between neighbour spins has been studied only by numerous Monte Carlo simulations (see e.g., reviews in [42, 41]). Unfortunately, we are unware of any studies of such a model when interactions decay with distance according to equation (1.1).
For systems with quenched disorder, the thermodynamics of physical observables can be analysed from the configurational average of the free energy [43]:
@@
$\overline{\left.\dots\right.}=\prod_{\bf x}\sum_{c_{\bf x}}p(c_{\bf x})(\dots)\,.$ (2.5)
To avoid averaging of the logarithm in equation (2.3) one can utilize the replica trick [44, 45, 46]
$\ln Z(\{c\})=\lim_{m\to 0}\frac{\big[Z(\{c\})\big]^{m}-1}{m}\,.$ (2.6)
The $m$th power of the configuration-dependent partition function is represented in the form of a functional integral using the Stratonovich–Hubbard transformation. This transformation introduces field variables $\vec{\phi}^{\alpha}_{{\bf x}}$, ($\alpha=\{1,\dots,m\}$), conjugated to the spin variables. Subsequently, the trace over the spin subsystem is taken, and after configurational averaging, the free energy is obtained (see e.g., [47]):
${F}\sim\int({\rm d}\vec{\phi})\mathrm{e}^{-{\cal H}_{\rm eff}},\qquad\int({\rm d}{\phi})=\prod_{\alpha=1}^{n}\prod_{\bf{x}}\int_{-\infty}^{\infty}{\rm d}{\phi}^{\alpha}_{\bf x}\,.$ (2.7)
@@
${\mathcal{H}}_{\rm eff}{=}\int\mathrm{d}^{d}x\left\{\frac{1}{2}\left[\tau|{\phi}^{\alpha}|^{2}{+}|\nabla^{\sigma/2}{\phi}^{\alpha}|^{2}\right]{+}\frac{u_{0}}{4!}\sum_{\alpha=1}^{m}|{\phi}^{\alpha}|^{4}{+}\frac{v_{0}}{4!}\bigg[\sum_{\alpha=1}^{m}({\phi}^{\alpha})^{2}\bigg]^{2}\right\}\,.$ (2.8)
In the limit $m\rightarrow 0$, the field theory (2.8) describes the critical properties of a LR interacting random Ising model. The parameter $\tau$ is connected with the temperature distance to the critical point. The bare coupling $u_{0}$ is proportional to the concentration $c$ and is positive whereas the bare coupling $v_{0}$ is proportional to c(c1)c(c-1) and thus is negative. The last term in (2.8) is present only for non-zero dilution: it is directly responsible for the fluctuations effective interaction due to the presence of impurities. We remind that (2.8) was obtained with the assumption of a weak quenched dilution (far above the percolation threshold). Note that theory (2.8) is valid only for $\sigma<2-\eta_{\rm SR}$. Since we are interested only in the long-distance properties of the theory, in the limit of a small wave vector $k\to 0$, we kept only the first non-trivial term of the expansion of the Fourier image of interactions $k^{\sigma}$.
@@ 3 RG approach
Within the field-theoretical RG approach [48, 49, 50], the analysis of the critical properties of $\phi^{4}$ models is based on the calculation of the so-called one-particle irreducible vertex functions. These functions are perturbatively calculated around the Gaussian model [$u=v=0$ in equation (2.8)] in the coupling constants $u$ and $v$. However, these calculations result in integrals that are divergent at the upper critical dimension, which in our case is $d_{u}=2\sigma$. To remove these divergencies, a controlled rearrangement of the perturbative series for the vertex functions is performed by imposing certain normalizing conditions. In what follows, we use the massless minimal subtraction scheme [51, 52]. We express bare parameters via the renormalized ones, introducing the renormalizing factors of the field $\phi$, $Z_{\phi}$, the couplings, $Z_{u}$ and $Z_{v}$ and for operator $\phi^{2}$, $Z_{\phi^{2}}$ ensuring finiteness of two-point vertex function with one $\phi^{2}$ insertion.
Then, the scaling bahaviour can be analysed from RG $\beta$- and $\gamma$-functions characterizing the change of the vertex functions under the RG transformation. They are introduced as follows:
$\beta_{u}=\frac{\partial u}{\partial\ln{\mu}},\quad\beta_{v}=\frac{\partial v}{\partial\ln{\mu}},\quad\gamma_{\phi}=\frac{\partial Z_{\phi}}{\partial\ln{\mu}},\quad\overline{\gamma}_{\phi^{2}}=-\frac{\partial\overline{Z}_{\phi^{2}}}{\partial\ln{\mu}},$ (3.1)
where $\overline{Z}_{\phi^{2}}=Z_{\phi^{2}}Z_{\phi}$ and the derivatives are taken at fixed bare parameters. Here, $\mu$} is the scale parameter introduced in the same way as in [19].
@@
The stability of the FPs (3.2) is defined by the stability matrix $\left(\begin{smallmatrix}{\partial\beta_{{u}}}/{\partial u}&{\partial\beta_{{u}}}/{\partial v}\\ {\partial\beta_{{v}}}/{\partial u}&{\partial\beta_{{v}}}/{\partial v}\end{smallmatrix}\right)$. If this matrix, evaluated at a FP $(u^{*},v^{*})$, has both its eigenvalues $\lambda_{1}$ and $\lambda_{2}$ with positive real parts, then this FP is stable.
@@
All other critical exponents can be calculated from the scaling relations [48, 49, 50].
We work in the three-loop approximation, using the results of loop integral calculations [19, 24]. Note that our procedure differs. The authors of references [19, 24] fixed spatial dimension $d$ and looked for poles in the deviation of the decay parameter from the real space dimension $\sigma=(d-\epsilon)/2$, while we fix $\sigma$ and look for poles in the deviations of spatial dimension from the upper critical dimension $d=2\sigma-\epsilon$. The result of our calculations of the RG functions (3.1) within the three-loop approximation is as follows:
$\beta_{u}$ $=$ $u\Big[-\epsilon+\frac{3u}{2}+2v-\frac{1}{18}\left(27u^{2}+72uv+42v^{2}\right)D_{\sigma}$ (3.4)
@@
${}+\frac{1}{216}(3u+2v)\left(9u^{2}+30uv+28v^{2}\right)H_{\sigma}\Big],$
$\beta_{v}$ $=$ $v\Big[-\epsilon+6u+\frac{4}{3}v-\frac{1}{24}\left(9u^{2}+36uv+22v^{2}\right)D_{\sigma}$ (3.5)
${}+\frac{1}{24}\left(21u^{3}+94u^{2}v+129uv^{2}+48v^{3}\right)D_{\sigma}^{2}+\frac{1}{108}v\left(63u^{2}+111uv+44v^{2}\right)E_{\sigma}$
${}+\frac{1}{{72}}\left(9u^{3}+30u^{2}v+45uv^{2}+16v^{3}\right)F_{\sigma}+\frac{1}{108}(3u+4v)\left(3u^{2}+6uv+2v^{2}\right)G_{\sigma}$
${}+\frac{1}{108}v\left(9u^{2}+24uv+11v^{2}\right)H_{\sigma}\Big],$
$\bar{\gamma}_{\phi^{2}}$ $=$ $\frac{1}{2}u+\frac{1}{3}v-\frac{1}{12}\left(3u^{2}+6uv+2v^{2}\right)D_{\sigma}$ (3.6)
${}+\frac{1}{864}\left(378u^{3}+1125u^{2}v+990uv^{2}+220v^{3}\right)D_{\sigma}^{2}+\frac{1}{432}v\left(9u^{2}+18uv+4v^{2}\right)E_{\sigma}$
${}+\frac{1}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}96}}\left(6u^{3}+19u^{2}v+18uv^{2}+4v^{3}\right)F_{\sigma}+\frac{1}{216}(3u+2v)\left(3u^{2}+6uv+2v^{2}\right)G_{\sigma},$